The attractive electrostatic force between the point charges 8.11×10−6 C and Q has a magnitude of 0.495 N when the separation between the charges is 5.40 m.

find the sign and magnitude of the charge Q
Answer #1

From Coulomb's inverse square law the force of attraction or repulsion is given by F= \frac{1}{4\pi \epsilon _{0}}\frac{Q_{1}Q_{2}}{R^{2}}

Here in our problem 0.495 = 9x109x8.22x10-6 xQ/(5.4)2

Therefore the charge on the second body is

                                 Q = 0.495x(5.4)2/ (9x109x8.11x10-6)

                                  Q = 198x10-3C

Since the force is attractive the charges should be opposite in charge. Hence this charge should be negative.

                                Therefore Q=-198x10-3C